# A Modular, Adaptive, and Scalable Quantum Factoring Algorithm

### v1

@@ Abstract
Shor’s algorithm for integer factorization offers an exponential speedup over classical methods but remains impractical on Noisy Intermediate-Scale Quantum (NISQ) hardware due to the need for many coherent qubits and very deep circuits. Building on our recent work on adaptive and windowed phase-estimation methods, we have developed a modular, windowed formulation of Shor’s algorithm that mitigates these limitations by restructuring phase estimation into shallow, independent circuit blocks that can be executed sequentially or in parallel, followed by lightweight classical postprocessing. This approach allows for a reduction in the size of the phase (or counting) register from thousands of qubits down to a small, fixed block size of only a few qubits (for example, three or four), while leaving the work register requirement unchanged. The independence of the blocks allows for parallel execution and makes the approach more compatible with near-term hardware than the standard Shor’s formulation. An additional feature of the framework is the overlap mechanism, which introduces redundancy between blocks and enables robust reconstruction of phase information, though zero-overlap configurations can also succeed in certain regimes. Numerical simulations verify the correctness of the modular formulation while also showing substantial reductions in counting qubits per block.
@@ 1 Introduction
Despite its theoretical power, the practical realization of Shor’s algorithm faces substantial barriers. In its standard form, the algorithm requires a large phase (or counting) register of $2n+1$ qubits, in addition to an $n$-qubit work register, together with long coherent circuits to implement repeated controlled modular exponentiation. Such resource demands exceed the capabilities of today’s Noisy Intermediate-Scale Quantum (NISQ) devices, which are constrained by limited qubit counts, short coherence times, and noise. A range of strategies have been explored to mitigate these requirements, including circuit-level optimizations techniques [3, 4, 5, 6].
@@
The remainder of the paper is organized as follows. Section 2 describes our proposed modular approach to integer factorization based on Shor’s algorithm. Section 3 presents numerical demonstrations of factoring small integers using shallow, quantum circuits. Computational complexity is discussed in Section 4. Section 5 concludes the paper with a summary of the results.
@@ 2.2 Carry-Aware Integration and Stitching
Further, carry errors can propagate across the boundaries of overlapping blocks and must be accounted for while stitching the full phase string from the individual block estimates. To resolve this, Algorithm 3 provides a robust, carry-aware stitching routine. The algorithm integrates the blocks from right to left, a process that naturally accounts for the direction of carry propagation. For any two consecutive blocks, a left candidate $s_{\text{left}}$ from block $\ell$ and a right candidate $s_{\text{right}}$ from block $\ell+1$, a consistency check is performed. The check uses the overlap size, $t_{\ell+1}$, to compare the tail of the left bitstring with the head of the right bitstring, while also accounting for a potential carry bit, $c$, from the right block. This is captured by the modulo arithmetic check: $(\mathrm{int}(\mathrm{tail}(s_{\ell},t_{\ell+1}),2)-c)\pmod{2^{t_{\ell+1}}}=\mathrm{int}(\mathrm{head}(s_{\ell+1},t_{\ell+1}),2)$ (ref. Remarks 1 and 2). If the consistency check passes, the blocks are stitched together, removing the overlap. This iterative procedure is performed for all blocks, ultimately producing a set of complete, stitched candidates, $\mathcal{C}_{\text{stitched}}$.
@@ 2.3 Period Recovery and Factorization
Summary: Registers and controlled-unitary exponents use LSB indexing; measured bitstrings are MSB-first; integer conversion treats $b$ as MSB-first (no reversal); concatenation/slicing are MSB-first.
@@
1 Initialize an empty list $\mathcal{S}$ of candidate sets;
@@ 2 Initialize exponent_offset $\leftarrow 0$;
/* 1. Run a QPE block for each block of phase qubits */
@@ 6   Append $\mathcal{S}_{i}$ to $\mathcal{S}$;
9
/* 2. Integrate blocks and stitch candidates */
@@
/* 3. Recover period and factor N from the final candidates */
@@ 11 $(r,\text{factor})\leftarrow\text{RecoverPeriodAndFactor}(\mathcal{C}_{\text{stitched}},a,N)$;
14
15return “No non-trivial factor found”;
@@
/* 1. Construct and execute the quantum circuit */
@@ 8 Measure the counting register in the computational basis to obtain a bitstring $b$;
/* 2. Process measurement outcomes */
@@ 9 Execute the circuit for shots total runs, recording the frequency of each outcome;
/* 3. Select and return candidate phases */
@@ 10 Select the top num_selected_candidates bitstrings based on their frequencies;
/* 1. Integrate blocks from right to left */
@@ 17 $\mathcal{R}\leftarrow\mathcal{R}_{\text{new}}$;
20
/* 2. Stitch the final sequences */
@@ 23   Compute $n_{\text{total}}\leftarrow\sum_{i=1}^{B}(m_{i}-t_{i})$;
27
28return $\mathcal{C}_{\text{stitched}}$;
@@
Output : A tuple $(r,\text{factor})$ where factor is a non-trivial factor of NN, or $(\texttt{None},\texttt{None})$ if no factor is found.
/* Per-Candidate Period Estimates via Continued Fractions */
@@ 11          $f_{2}\leftarrow\text{gcd}(a^{r/2}+1,N)$;
22return $(\texttt{None},\texttt{None})$;
@@ 3 Numerical Examples
To illustrate our proposed modular approach to Shor’s algorithm (Algorithm 1), we present several numerical examples. These small-scale demonstrations verify the correctness of the approach and highlight how varying block sizes and overlaps affect the phase-estimation procedure, resource requirements, and the subsequent recovery of non-trivial factors. We begin with the case $N=15$, and then consider alternative block configurations and a larger composite, $N=221$.
@@ 3.2 Numerical Example: Quantum Factoring of $N=15$ with Fewer Qubits
The QPE is executed in two independent blocks, denoted as $B=2$. The block sizes are defined by the vector $\mathbf{m}=[3,4]$, while the corresponding overlaps are given by $\mathbf{t}=[0,2]$. For each block, a circuit is run to measure the most significant qubits of the phase register. Table 2 summarizes the most frequent outcomes, with probabilities derived empirically from simulation shots.
@@ 3.3 Numerical Example: Quantum Factoring of $N=15$ with Zero Overlap
The QPE is executed in two sequential blocks, denoted as $B=2$. The block sizes are defined by the vector $\mathbf{m}=[3,3]$, while the corresponding overlaps are given by $\mathbf{t}=[0,0]$ (i.e. the blocks do not overlap). For each block, a circuit is run to measure the most significant qubits of the phase register. Table 3 summarizes the most frequent outcomes, with probabilities derived empirically from simulation shots.
@@
3.4 Numerical Example: Quantum Factoring of $N=221$ with Fewer Qubits
@@
The QPE is executed in four sequential blocks, denoted as $B=4$. The block sizes are defined by the vector $\mathbf{m}=[3,3,4,3]$, while the corresponding overlaps are given by $\mathbf{t}=[0,2,2,2]$. For each block, a circuit is run to measure the most significant qubits of the phase register. Table 4 summarizes the most frequent outcomes, with probabilities derived empirically from simulation shots.

### v5

@@ Abstract
Shor’s algorithm for integer factorization offers an exponential speedup over classical methods but remains impractical on Noisy Intermediate-Scale Quantum (NISQ) hardware due to the need for many coherent qubits and very deep circuits. Building on our recent work on adaptive and windowed phase-estimation methods, we have developed a modular, windowed formulation of Shor’s algorithm that potentially mitigates these limitations by restructuring phase estimation into shallow, independent circuit blocks that can be executed sequentially or in parallel, followed by lightweight classical postprocessing. This approach allows for a reduction in the size of the phase (or counting) register from a large number of qubits down to a small, fixed block size of only a few qubits (for example, three or four phase qubits were sufficient for the computational examples considered in this work), while leaving the work register requirement unchanged. The independence of the blocks allows for parallel execution and makes the approach more compatible with near-term hardware than the standard Shor’s formulation. An additional feature of the framework is the overlap mechanism, which introduces redundancy between blocks and enables robust reconstruction of phase information, though zero-overlap configurations can also succeed in certain regimes. Numerical simulations verify the correctness of the modular formulation while also showing substantial reductions in counting qubits per block.
@@ 1 Introduction
Despite its theoretical power, the practical realization of Shor’s algorithm faces substantial barriers. In its standard form, the algorithm requires a large phase (or counting) register of $2n+1$ qubits, in addition to an $n$-qubit work register, together with long coherent circuits to implement repeated controlled modular exponentiation to factor an $n$-bit integer. Such resource demands exceed the capabilities of today’s Noisy Intermediate-Scale Quantum (NISQ) devices, which are constrained by limited qubit counts, short coherence times, and noise. A range of strategies have been explored to mitigate these requirements, including circuit-level optimizations techniques [3, 4, 5, 6].
@@
The remainder of the paper is organized as follows. Section 2 describes our proposed modular approach to integer factorization based on Shor’s algorithm. Section 3 presents numerical demonstrations of integer factorization using shallow, quantum circuits. A deterministic candidate-based success metric is presented in Section 5. Computational complexity is discussed in Section 6. Section 7 concludes the paper with a summary of the results.
@@ 2.2 Carry-Aware Integration and Stitching
Further, carry errors can propagate across the boundaries of overlapping blocks and must be accounted for while stitching the full phase string from the individual block estimates. To resolve this, Algorithm 3 provides a robust, carry-aware stitching routine. The algorithm integrates the blocks from right to left, a process that naturally accounts for the direction of carry propagation. For any two consecutive blocks, a left candidate $s_{\text{left}}$ from block $\ell$ and a right candidate $s_{\text{right}}$ from block $\ell+1$, a consistency check is performed. The check uses the overlap size, $t_{\ell+1}$, to compare the tail of the left bitstring with the head of the right bitstring, while also accounting for a potential carry bit, $c$, from the right block. This is captured by the modulo arithmetic check: $(\mathrm{int}(\mathrm{tail}(s_{\ell},t_{\ell+1}),2)-c)\pmod{2^{t_{\ell+1}}}=\mathrm{int}(\mathrm{head}(s_{\ell+1},t_{\ell+1}),2)$ (ref. Remarks 1 and 3). If the consistency check passes, the blocks are stitched together, removing the overlap. This iterative procedure is performed for all blocks, ultimately producing a set of complete, stitched candidates, $\mathcal{C}_{\text{stitched}}$.
@@ 2.3 Period Recovery and Factorization
To summarize, registers and controlled-unitary exponents use LSB indexing; measured bitstrings are MSB-first; integer conversion treats $b$ as MSB-first (no reversal); concatenation/slicing are MSB-first.
Remark 2.
[…]
where $\{k_{j}\}_{j=0}^{L-1}$ is a set of non-negative integer exponents chosen from a suitable range, and $\{w_{j}\}_{j=0}^{L-1}$ are the complex weights (amplitudes) satisfying the normalization condition $\sum_{j=0}^{L-1}|w_{j}|^{2}=1$.
This construction yields a general multi-term superposition in the target register. The exponents $k_{j}$ can be chosen in various ways, for instance, in an arithmetic progression where the exponents increase linearly (e.g., $k_{j}=A+B\cdot j$), or in a non-linear progression where they increase quadratically, exponentially, or according to any other desired non-linear pattern. The goal is to create a non-uniform initial target state such that one phase dominates in the superposition, simplifying subsequent processing.
[…]
where $\kappa_{0}=\texttt{exponent\_offset}$ (ref.  Algorithm 2).
[…]
@@
1
21exInitialize an empty list $\mathcal{S}$ of candidate sets;
@@ 3 Initialize exponent_offset $\leftarrow 0$;
4
1ex/* 1. Run a QPE block for each block of phase qubits */
@@ 8   Append $\mathcal{S}_{i}$ to $\mathcal{S}$;
11
12
1ex/* 2. Integrate blocks and stitch candidates */
@@
14
1ex/* 3. Recover period and factor N from the final candidates */
@@ 15 $(r,\text{factor})\leftarrow\text{RecoverPeriodAndFactor}(\mathcal{C}_{\text{stitched}},a,N)$;
18
19
201exreturn “No non-trivial factor found”;
@@
1
1ex/* 1. Construct and execute the quantum circuit */
@@ 10 Measure the counting register in the computational basis to obtain a bitstring $b$;
11
1ex/* 2. Process measurement outcomes */
@@ 12 Execute the circuit for shots total runs, recording the frequency of each outcome;
13
1ex/* 3. Select and return candidate phases */
@@ 14 Select the top num_selected_candidates bitstrings based on their frequencies;
1
1ex/* 1. Integrate blocks from right to left */
@@ 18 $\mathcal{R}\leftarrow\mathcal{R}_{\text{new}}$;
21
22
1ex/* 2. Stitch the final sequences */
@@ 25   Compute $n_{\text{total}}\leftarrow\sum_{i=1}^{B}(m_{i}-t_{i})$;
29
30
311exreturn $\mathcal{C}_{\text{stitched}}$;
@@
Output : A tuple $(r,\text{factor})$ where factor is a non-trivial factor of $N$, or $(\texttt{None},\texttt{None})$ if no factor is found.
1
1ex/* Per-Candidate Period Estimates via Continued Fractions */
@@ 12          $f_{2}\leftarrow\text{gcd}(a^{r/2}+1,N)$;
22
23
241exreturn $(\texttt{None},\texttt{None})$;
@@
2.4 Rationale for Using a Non-Standard Initial Target State
[…]
To enhance robustness and spectral diversity in our adaptive and windowed QPE framework, as discussed in Remark 2, we prepare a more general superposition in the target register:
[…]
where $\{k_{j}\}_{j=0}^{L-1}$ is a set of appropriately chosen non-negative integer exponents. The exponents $\{k_{j}\}$ and amplitudes $\{w_{j}\}$ can be selected to produce a non-uniform distribution over eigenstates, resulting in one or more dominant eigencomponents. This, in turn, can improve the distinguishability of valid phase estimates.
[…]
Example ($L=2$). For a simple two-term initialization
[…]
we have $w_{0}=w_{1}=1/\sqrt{2}$ and $w_{s\geq 2}=0$. Using the above formulation, the decomposition becomes
[…]
yielding spectral weights
$|c_{k}|^{2}=\frac{2}{r}\cos^{2}\!\left(\frac{\pi k}{r}\right).$
Thus, the modified initialization redistributes amplitude non-uniformly across the eigenstates of $U_{a}$, producing a distinct, structured pattern in the observed phase histogram. This structured modulation remains robust (across blocks) and enhances correlation between blocks (with overlap) and supports partial-block stitching under limited resolution, making it particularly advantageous in our windowed QPE approach.
Remark 4.
For simplicity num_selected_candidates is taken to be identical for all blocks, though in general each block may use its own value, for example $N_{i}^{\mathrm{sel}}$ for the $i$-th block.
Remark 5.
Further improvements to robustness in the overall order-finding process can be achieved by combining multiple phase candidates arising from different initialization choices or independent runs using the least common multiple (LCM) of their corresponding period estimates. Since the initialization in Equation 2 affects only the probability distribution and not the eigenphases themselves, the valid period $r$ remains consistent across such runs. Taking the LCM of independently obtained candidate periods amplifies stability and helps suppress spurious estimates introduced by measurement noise or imperfect block stitching.
@@ 3 Numerical Examples
To illustrate our proposed modular approach to Shor’s algorithm (Algorithm 1), we present several numerical examples. These small-scale demonstrations verify the correctness of the approach and highlight how varying block sizes and overlaps affect the phase-estimation procedure, resource requirements, and the subsequent recovery of non-trivial factors.
[…]
@@ 3.2 Numerical Example: Quantum Factoring of $N=15$ with Fewer Qubits
The QPE is executed in two independent blocks, denoted as $B=2$. The block sizes are defined by the vector $\mathbf{m}=[3,4]$, while the corresponding overlaps are given by $\mathcal{T}=[0,2]$. For each block, a circuit is run to measure the most significant qubits of the phase register. Table 2 summarizes the most frequent outcomes, with probabilities derived empirically from simulation shots.
@@ 3.3 Numerical Example: Quantum Factoring of $N=15$ with Zero Overlap
The QPE is executed in two sequential blocks, denoted as $B=2$. The block sizes are defined by the vector $\mathbf{m}=[3,3]$, while the corresponding overlaps are given by $\mathcal{T}=[0,0]$ (i.e. the blocks do not overlap). For each block, a circuit is run to measure the most significant qubits of the phase register. Table 3 summarizes the most frequent outcomes, with probabilities derived empirically from simulation shots.
@@
3.4 Numerical Example: Quantum Factoring of $N=221$
@@
The QPE is executed in four sequential blocks, denoted as $B=4$. The block sizes are defined by the vector $\mathbf{m}=[3,3,4,3]$, while the corresponding overlaps are given by $\mathcal{T}=[0,2,2,2]$. For each block, a circuit is run to measure the most significant qubits of the phase register. Table 4 summarizes the most frequent outcomes, with probabilities derived empirically from simulation shots.